Top rated cars is a world manufacture of cars and trucks. They maintain a reserve on the balance sheet for future warranty claims. The balance in this account at December 31,2015,2016, 2017, and 2018 ( in Millions) was $315.0, $329.0, $354.7 and $361.9 respectively. In addition, revenues  for the same periods wer ( in millions). $10,000.8, $ 10,122.0, $ 10,098.6. Form this data, one could surmise that the quality of their cars and trucks is improving. True or False
Answer #1

Year Revenues Warranty Claims

2015 10,000.08 315.0

2016 10,122.00 329.0

2017 10,098.60 354.7

2018 Not Given 361.9

the change in reserve balance for future warranty claims.

1. (329 - 315)/315 =4.4%

2. (354-329)/329 =7.8%

3.(361-354.7)/354.7 = 2.0%

From the above change can say that from the year 2015 - 2016 taking as a base for stating the quality Improvment.

2016-2017 = Quality of cars and truck was going down, as provision for warranty claims increasing, but the revenue is declining.

From - 2017-2018 = Quality is again considered in focus and gradually started Improving but due to non-availability of revenue fig of 2018. we can not compare change in reserve balance with revenue.

Conclusion - from the decline in the change in the reserve balance of warranty claim - Quality of cars and trucks are gradually Improving.

True.